Circular No. 6717

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1997 P2 T. Gehrels, Lunar and Planetary Laboratory, reports the automatic

discovery of a comet by Spacewatch. Available CCD observations:

1997 UT α_{2000} δ_{2000} m_{1} Observer
Aug. 12.24372 21^{h}40^{m}04.^{s}76 −6^{o}37′47′′.8 18.6 Gehrels
12.26467 21 40 04.28 −6 37 45.0 19.0  _{”}
12.28579 21 40 03.67 −6 37 44.7 19.0  _{”}
13.33624 21 39 36.76 −6 36 39.3  Balam
13.35463 21 39 36.28 −6 36 37.1  _{”}
13.36181 21 39 36.08 −6 36 36.8  _{”}

T. Gehrels (Kitt Peak). 0.9-m Spacewatch telescope. Tail 1′ long in p.a.

325^{o}. D. D. Balam (Dominion Astrophysical Observatory). 1.8-m Plaskett reflec-

tor. Clearly cometary, quite condensed, no tail observed.

SUPERNOVA 1997dc IN NGC 7678 A. Piemonte, S. Benetti and M. Turatto, European Southern Observa-

tory, observed SN 1997dc with the 1.54-m Danish telescope (+ DFOSC) on Aug. 11.2 UT at La Silla: “The spectrum (range 400-950 nm, resolution 1.1 nm) shows He I lines with P-Cyg profiles at 587.6 (possibly contaminated by Na D), 667.8 and 706.5 nm with an expansion velocity of 7500 km/s. Similar expansion velocity is shown by the Si II doublet 635.5 nm, while the expansion velocity measured from the Ca II infrared triplet is 15 000 km/s. The strong He I lines and the relatively low expansion velocity are consistent with a type Ib classification around maximum light. The spectrum shows a red continuum indicative of strong reddening, as suggested by the intense interstellar Na D absorption lines. The redshift at the location of the supernova is z = 0.012, as measured from the Na D absorption and the emission lines of an underlying H II region.”

COMET 2P/ENCKE Visual m_{1} estimates: June 26.46 UT, 7.8 (A. Pearce, City Beach, W.

Australia, 20 × 80 binoculars); 30.46, 8.4 (Pearce); July 3.35, 8.5 (D. A. J. Seargent, The Entrance, N.S.W., 25 × 100 binoculars); 7.45, 9.7 (Pearce, Subiaco, W. Australia, 0.2-m reflector). CCD m_{2} estimates by G. J. Garradd, Loomberah, N.S.W. (0.25-m reflector): July 7.40 UT, 16.6; 10.47, 16.1; 14.44, 16.0; 21.37, 16.2; 24.49, 16.9; 28.50, 17.1; 30.47, 17.4.

1997 August 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden